OpenStudy (anonymous):

solve for theta? -sin2theta=0

OpenStudy (anonymous):

\[\theta =0 \]

OpenStudy (anonymous):

or If you want a better answer \[\theta=n\pi\] \[n \in 0,1,2,...\]

OpenStudy (anonymous):

\[\theta=\frac{n\pi}{2}\]Sorry, I forgot it's \(2\theta\)